Carrie is three times older than Julio. If the sum of their ages is 48, how old is Julio

Answer by Alan3354(69443)

Carrie is three times older than Julio. If the sum of their ages is 48, how old is Julio
----------------
J = Julio's age
C = Carrie's age
===================
C = J + 3J = 4J
J + 4J = 48
J = 9.6
C = 28.8
==============================
1 time older = 2x as old
2 times older = 3x as old
etc.
======================================
People who don't know what they're doing should not try to write math books.